The decay of a radioisotope has a rate constant of 0.09 s-1 . How much time is required for 3/4 of a 1-gram sample of the isotope to decay?

Show your work.
Answer #1

we have:

[A]o = 1 g

3/4 of 1 g has decayed.

[A] = 1/4 of 1 g = 0.25 g

k = 0.09 s-1

use integrated rate law for 1st order reaction

ln[A] = ln[A]o - k*t

ln(0.25) = ln(1) - 0.09*t

-1.386 = 0 - 0.09*t

0.09*t = 1.386

t = 15.4 s

Answer: 15.4 s